# Zero-Day Audio DeepFake Detection via Retrieval Augmentation and Profile Matching

### v2

@@ Abstract
Modern audio deepfake detectors built on foundation models and large training datasets achieve promising detection performance. However, they struggle with zero-day attacks, where the audio samples are generated by novel synthesis methods that models have not seen from reigning training data. Conventional approaches fine-tune the detector, which can be problematic when prompt response is needed. This paper proposes a training-free retrieval-augmented framework for zero-day audio deepfake detection that leverages knowledge representations and voice profile matching. Within this framework, we propose simple yet effective retrieval and ensemble methods that reach performance comparable to supervised baselines and their fine-tuned counterparts on the DeepFake-Eval-2024 benchmark, without any additional model training. We also conduct ablation on voice profile attributes, and demonstrate the cross-database generalizability of the framework with introducing simple and training-free fusion strategies.
@@ III-A Self-supervised learning models
Profile extractor. We employ vox_profile [^4] https://github.com/tiantiaf0627/vox-profile-release [6] to extract profile feature vectors using well-trained foundation models. We construct the features by concatenating four attributes: 1) Age: a scalar representing age value in terms of 10 uniformed intervals from 0 to 100 years old555e.g., A value of 7 means an estimated age of 61-70 years old.; 2) Gender: a 2-dimensional one-hot vector for male/female classification; 3) Emotion: a scalar indicating the emotion trait (anger, fear, neutral, etc.) and 256-dimensional embedding vector extracted from the intermediate of the foundational model; 4) Voice quality: a 25-dimensional vector capturing five perceptual aspects: pitch, volume, clarity, rhythm, and voice texture, following the taxonomy defined in [5]. More details about how the representations are extracted can be found in [6].
@@ VI Conclusion
In this study, we have addressed zero-day attacks in audio deepfake detection, where systems must use existing knowledge to rapidly handle unseen attacks. We have introduced an retrieval augmentation framework based on knowledge databases and various related strategies, without any manner of fine-tuning or additional training. The framework uses both state-of-the-art CM and voice profile representations. Leveraging DE2024 with various duration conditions, we have found that the proposed method, especially when combining SSL and profile features from the retrieved data, achieves competitive performance. Finally, we have evaluated the framework on the cross-database scenario and show that the even training-free approach leveraging knowledge databases may be affected by domain mismatch. Future work may explore acquiring temporal information via frame-level SSL features, modeling profile information more effectively, and extending the framework to multi-classification ADD-related tasks.

### v3

@@ Abstract
Modern audio deepfake detectors built on foundation models and large training datasets achieve promising detection performance. However, they struggle with zero-day attacks, where the audio samples are generated by novel synthesis methods that models have not seen from reigning training data. Conventional approaches fine-tune the detector, which can be problematic when prompt response is needed. This paper proposes a training-free retrieval-augmented framework for zero-day audio deepfake detection that leverages knowledge representations and voice profile matching. Within this framework, we propose simple yet effective retrieval and ensemble methods that reach performance comparable to supervised baselines and their fine-tuned counterparts on the DeepFake-Eval-2024 benchmark, without any additional model training. We also conduct ablation on voice profile attributes, and demonstrate the cross-database generalizability of the framework with introducing simple and training-free fusion strategies [^1] This study is supported by JST AIP Acceleration Research (JPMJCR24U3) and the New Energy and Industrial Technology Development Organization (NEDO, JPNP22007). This study was partially carried out using the TSUBAME4.0 supercomputer at the Institute of Science Tokyo..
@@ III-A Self-supervised learning models
Profile extractor. We employ vox_profile [^5] https://github.com/tiantiaf0627/vox-profile-release [6] to extract profile feature vectors using well-trained foundation models. We construct the features by concatenating four attributes: 1) Age: a scalar representing age value in terms of 10 uniformed intervals from 0 to 100 years old666e.g., A value of 7 means an estimated age of 61-70 years old.; 2) Gender: a 2-dimensional one-hot vector for male/female classification; 3) Emotion: a scalar indicating the emotion trait (anger, fear, neutral, etc.) and 256-dimensional embedding vector extracted from the intermediate of the foundational model; 4) Voice quality: a 25-dimensional vector capturing five perceptual aspects: pitch, volume, clarity, rhythm, and voice texture, following the taxonomy defined in [5]. More details about how the representations are extracted can be found in [6].
@@ VI Conclusion
In this study, we have addressed zero-day attacks in audio deepfake detection, where systems must use existing knowledge to rapidly handle unseen attacks. We have introduced an retrieval augmentation framework based on knowledge databases and various related strategies, without any manner of fine-tuning or additional training. The framework uses both state-of-the-art CM and voice profile representations. Leveraging DE2024 with various duration conditions, we have found that the proposed method, especially when combining SSL and profile features from the retrieved data, achieves competitive performance.
Finally, we have evaluated the framework on the cross-database scenario and show that the even training-free approach leveraging knowledge databases may be affected by domain mismatch. Future work may explore acquiring temporal information via frame-level SSL features, modeling profile information more effectively, and extending the framework to multi-classification ADD-related tasks.